3 ladies &3 gents can be seated at a round table so that any two & only two of the ladies sit together.the number of ways is
Answer by Edwin McCravy(20055)

There are only these two GENERAL ways to seat them. [Rotations of
these two are not counted as separate GENERAL seating arrangements.] 

       L                          L
   L       G                  L       G  
   
   G       L                  G       G                 
       G                          L

Choose the GENERAL way to seat the 6 in 2 ways.
Arrange the gents in the three G seats in 3! SPECIFIC ways.
Arrange the ladies in the three L seats in 3! SPECIFIC ways.

Answer: 2*3!*3! = 2*6*6 = 72 ways. 

Edwin